OpenStudy (anonymous):

A student committee consists of 20 boys and 15 girls. A team of 8 students is selected randomly from the committee to participate in an activity and the team must consists of at least 3 boys and at least 3 girls. Find the probability that the teams can be formed if the team consists of 4 boys already.

OpenStudy (anonymous):

This will take me a few mins sorry im just thinking it through my head

OpenStudy (anonymous):

THe probability that they can be or will be?

OpenStudy (anonymous):

can be

OpenStudy (anonymous):

Well with 15 Girls left 4 slots on the team open and 16 boys left i would think 100%

OpenStudy (anonymous):

There seems to be no rules whether or not they can join a team

OpenStudy (anonymous):

" the team must consists of at least 3 boys and at least 3 girls."

OpenStudy (anonymous):

ATLEAST not ONLY so it would be 100% wouldnt it?

OpenStudy (anonymous):

If there are 7 boys, only 1 girl, then the team cannot be formed. So, it will NOT be 100%

OpenStudy (anonymous):

There arent multiple teams so there wouldnt be a number limit on how many of each we can add

OpenStudy (anonymous):

I suggest you to read the question carefully again.

OpenStudy (anonymous):

IT DOES NOT SAY IT HAS TO HAVE 3 MORE BOYS AND GIRLS ATLEAST MEANING THAT 4 BOYS AND 4 GIRLS WOULD WORK

OpenStudy (anonymous):

4 boys 4 girls is just one way It can be 3 boys 1 girl, then it fails.

OpenStudy (anonymous):

Oh i am sorry it is written in such a way that it was slightly dumbfounding me how my way wouldnt work

OpenStudy (anonymous):

But since there are 4 boys already there could be 4/4 5/3 6/2 7/1 8/0 so the probability would be 40% or 2/5

OpenStudy (anonymous):

Im sorry

OpenStudy (anonymous):

No...

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Your assumption in your calculation is that boys and girls will be picked with equal probability but they are not. There are more boys than girls.

OpenStudy (anonymous):

@kewlgeek555 Can you help me with this kewl?

OpenStudy (anonymous):

And for now i will use 40% as my base and will edit that to get the correct percent im sorry for not doing too well i stopped taking statistics last year but im pretty good compared to most people when it comes to problems that have an equation

OpenStudy (anonymous):

Oh, maybe you can help me with my other problem then. Let me solve the problem first though. :)

OpenStudy (anonymous):

Ok so selecting a girl has a 75% chance compared to selecting a boy not the full chance.

OpenStudy (anonymous):

GTG Be back in a bit

OpenStudy (anonymous):

LOL I think I got it.

OpenStudy (anonymous):

Let B = boy(s), G = girl(s) \[P(4B) = \frac{16}{31}\times\frac{15}{30}\times\frac{14}{29}\times\frac{13}{28}\times\frac{4!}{4!}=\frac{52}{899}\]\[P(3B1G) = \frac{16}{31}\times\frac{15}{30}\times\frac{14}{29}\times\frac{15}{28}\times\frac{4!}{3!}=\frac{240}{899}\]\[P(2B2G) = \frac{16}{31}\times\frac{15}{30}\times\frac{15}{29}\times\frac{14}{28}\times\frac{4!}{2!2!}=\frac{360}{899}\]\[P(1B3G) = \frac{16}{31}\times\frac{15}{30}\times\frac{14}{29}\times\frac{13}{28}\times\frac{4!}{3!}=\frac{208}{899}\]\[P(4G) = \frac{15}{31}\times\frac{14}{30}\times\frac{13}{29}\times\frac{12}{28}\times\frac{4!}{4!}=\frac{39}{899}\] So, \[\text{P(can form a team) = }\frac{P(1B3G) + P(4G)}{P(4B)+P(3B1G)+P(2B2G)+P(1B3G) + P(4G)} = \frac{247}{899}\] Does this make sense?

OpenStudy (anonymous):

The last fraction after the equal sign in the second last line is \(\large\frac{247}{899}\).

OpenStudy (anonymous):

Ok good lol

OpenStudy (anonymous):

The other one?

OpenStudy (anonymous):

Haha! You want some challenges?

OpenStudy (anonymous):

Sure just not ones that cant be solved with a percentage or a fraction i prefer not to do large equations in my head (I do much worse when i try to work on paper) please

OpenStudy (anonymous):

Hmm... I do have something I cannot solve at the moment... Hold on.

OpenStudy (anonymous):

Please show me

OpenStudy (anonymous):

Yeah hes making a new one kewl